OpenStudy (anonymous):

what is the slope y=8x-6

OpenStudy (hoblos):

8

OpenStudy (anonymous):

thanks ask me a question too see if i can answer it

OpenStudy (anonymous):

linear equation is y = mx + b wher m is the slope therfore 8 is the slope

OpenStudy (hoblos):

welcome :D what is the slope of 2y=-6x+4

OpenStudy (anonymous):

im thinking negative 6???

OpenStudy (hoblos):

no!! the slope is equal to the coefficient of x only if the coefficient of y is y so here you must divide by 2 first to get y=-3x+2 then the slope is -3

OpenStudy (hoblos):

coefficient of y is 1 sorry

OpenStudy (anonymous):

so the slope is 1

OpenStudy (mathmate):

"get y=-3x+2 then the slope is -3 "

OpenStudy (anonymous):

okay well can you give me a question like this 2y=-6x+4

OpenStudy (mathmate):

3y=12x+24

OpenStudy (anonymous):

okay so would i be dividing in this equagtion

OpenStudy (mathmate):

everything by the coefficient of y.

OpenStudy (anonymous):

okay so i would be div

OpenStudy (anonymous):

so i would be dividing by 3 right

OpenStudy (mathmate):

right, what do you get?

OpenStudy (anonymous):

so it would be 3??

OpenStudy (mathmate):

You would divide every term of the equation by 3 to get: y=4x+6 What is the slope?

OpenStudy (anonymous):

oh 4

OpenStudy (mathmate):

Good job! Remember, before you find the slope, the coefficient of y must be 1, alright?

OpenStudy (anonymous):

ohkay that makes sense

OpenStudy (mathmate):

Great!

OpenStudy (hoblos):

good job mathmate :D

OpenStudy (anonymous):

okay one more question give me a good one

OpenStudy (hoblos):

3x + 2y = 4 try this

OpenStudy (hoblos):

Qutira !!

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

3 right??

OpenStudy (hoblos):

nope!! think more and remember what i told you before ;)

OpenStudy (anonymous):

6??

OpenStudy (anonymous):

6

OpenStudy (hoblos):

also no i will make it easier to you it can be written as 2y =-3x+ 4 what is the slope then?

OpenStudy (anonymous):

am i multiplying

OpenStudy (hoblos):

you just have to divide by 2 to get y=-3/2x +2 => the slope is -3/2 get it?

OpenStudy (anonymous):

ehh alittle! but gtg

OpenStudy (hoblos):

lol bye :)) whenever you need help you can count on us 0k ;)